# Learning Attentive Neural Processes for Planning with Pushing Actions

### v1

@@ VI Conclusion
We present: the Pushing Neural Process, an Attentive Neural Process architecture to learn push outcome prediction of a novel object instance, and NPT-DPW, a new planner that integrates the PNP to produce good plans under reasonable computational runtime budgets.
Our initial experiments demonstrate promising capabilities; however, we acknowledge certain limitations that we aim to address in subsequent research. For instance, the PNP exhibited underfitting, which we attribute to the complexity and inherent instability of the simulation environment used to generate the training data. To address this, we hope to apply the NPT-DPW to a broader spectrum of tasks, including scenarios with simpler dynamics and transition models to yield clearer insights. Secondly, we observed occasional erratic behavior in generated plans, indicating opportunities for further model improvement. Addressing these issues may improve planning performance in both PFT-DPW and NPT-DPW. Finally, we hope to deploy the PNP on a physical robot platform in order to validate our approach more comprehensively and demonstrate applicability.

### v3

@@ VI Conclusion
We present: the Pushing Neural Process, an Attentive Neural Process architecture to learn push outcome prediction of a novel object instance, and NPT-DPW, a new planner that integrates the PNP to produce good plans under reasonable computational runtime budgets. Additionally, we demonstrated a method of learning a representation of an object’s properties, and how we can leverage that representation to maintain belief with a general use case planner.
Our initial experiments demonstrate promising capabilities; however, we acknowledge certain limitations that we aim to address in subsequent research. For instance, the PNP exhibited underfitting, which we attribute to the complexity and inherent instability of the simulation environment used to generate the training data. To address this, we hope to apply the NPT-DPW to a broader spectrum of tasks, including scenarios with simpler dynamics and transition models to yield clearer insights. Secondly, we observed occasional erratic behavior in generated plans, indicating opportunities for further model improvement. Addressing these issues may improve planning performance in both PFT-DPW and NPT-DPW. Finally, we hope to deploy the PNP on a physical robot platform to comprehensively validate our approach and demonstrate its applicability.